Solve each problem. Use your choice of methods.
How many liters (L) of a 10% silver iodide solution must be mixed with 7 L of a 4% silver iodide solution to get a 6% solution?

Answer by josmiceli(6785)

(.10*L + .04*7) / L + 7 = .06
.1*L + .28 = .06*L + .42
.04*L = .42 - .28
L = .14/.04
L = 3.5 liters of 10% solution
check
(.10*3.5 + .28) / (3.5 + 7) = .06
(.35 + .28) / 10.5 = .06
.63/10.5 = .06
OK